Find the point-slope of an equation of the line through the points at (5, -3) and (7, -4). Use (5, -3) as the point (x1, y1).
Thanks!

Answer by stanbon(75887)

Find the point-slope of an equation of the line through the points at (5, -3) and (7, -4). Use (5, -3)
The slope is [(-4--3)/(7-5)]
=-1/2
Pt-slope form is : y-y1=m(x-x1)
y-(-3)=(-1/2)(x-5)
y+3=(-1/2)x+(5/2)
y=(-1/2)x-(1/2)
Cheers,
Stan H.